OpenStudy (anonymous):

d^9/d^6 = (whole number)

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

give corresponding whole number

OpenStudy (anonymous):

reduce the left hand side

OpenStudy (anonymous):

are there any limitations on d?

OpenStudy (anonymous):

seems like it would be 1, because there is no whole number, only exp.

OpenStudy (anonymous):

just d^9\ d^6

OpenStudy (anonymous):

d^3 = whole number

OpenStudy (anonymous):

d could be any whole number or cube root or whole number times cube root

OpenStudy (anonymous):

so then wouldn't it just be 1, as d is understood as 1?

OpenStudy (anonymous):

why is d understood to be 1?

OpenStudy (anonymous):

just comparing to a linear equation? i don't know:(

geerky42 (geerky42):

\(\Huge d = \sqrt[3]{x}, x \in \mathbb{Z}\)

OpenStudy (anonymous):

\[x \in \mathbb{Z}^+\]

OpenStudy (anonymous):

whole number

OpenStudy (anonymous):

actually \[ x \in \mathbb{Z}^*\]

geerky42 (geerky42):

Or \(\large\mathbb{N} \)